1. Find the solution of the given initial value problem.
y prime - y = 3t(e^2t), y(0) = 1

y(t) = _____

2. Find the solution of the given initial value problem in explicit form.

y prime = (1-11x)y^2, y(0) = -1/6

Enclose numerators and denominators in parentheses. For example, (a-b) / (1+n).

y(x) = _____
1
Expert's answer
2020-06-11T20:04:49-0400

(i) Given

"\frac{dy}{dx} - y = 3te^{2t}" with "y(0) = 1"


Solving the differential equation,

Equation is linear in y and first order equation.


Integral factor will be "e^{ - \int dt} = e^{-t}"


multiplying this factor on both sides of equation,

"e^{-t}(\frac{dy}{dx} - y) = 3te^{2t}e^{-t}"

"e^{-t}(\frac{dy}{dx} - y) = 3te^{t}"

Now integration both sides with respect to x,

"e^{-t}y = \int 3te^{t}dt"

"e^{-t}y = 3[te^{t} - \int e^{t} dt]"

"e^{-t}y = 3[te^{t} -e^{t}]" "+ c"

"e^{-t}y = 3e^{t}(t-1) + c"

so "y = 3e^{2t}(t-1)+ce^{t}"

applying condition, "y(0) = 1,"


"1 = -3 +c\implies c=4"

so equation will be

"y = 3e^{2t}(t-1) + 4 e^t"



(ii) "\frac{dy}{dx} = (1-11x)y^2" with "y(0) = \frac{-1}{6}"

integrating both sides


"\int \frac{dy}{y^2} = \int (1-11x)dx"


"\frac{-1}{y} = x - \frac{11}{2}x^2 + c"


applying condition "y(0) = \frac{-1}{6}"

then "6 = 0 - 0 +c \implies c = 6"

then equation will be


"\frac{-1}{y} = x - \frac{11}{2}x^2 + 6"

"\frac{-1}{y} = \frac{2x - 11x^2 + 12}{2}"

then "{y} = \frac{-2}{2x - 11x^2 + 12}"